Let f be a differentiable function. Establish that identity |f'(z)|^2=(Ux)^2 + (Vx)^2 = (Uy)^2 + (Vy)^2
1
Expert's answer
2022-01-31T16:41:39-0500

Let f be a "\mathbb{C}"-differentiable function at a point "z\in\mathbb{C}". This means that there exist the following limit: "\lim\limits_{\Delta z\to 0}\frac{f(z+\Delta z)-f(z)}{\Delta z}=f'(z)".

In particular, if we restrict "\Delta z" to take only real values (in this case then "\Delta z=\Delta x"), then we conclude that there exists a partial derivative "\partial f/\partial x" at the point "z":

"f_x(z):=\frac{\partial f}{\partial x}(z):=\lim\limits_{\Delta x\to 0}\frac{f(z+\Delta x)-f(z)}{\Delta x}=f'(z)"

In the similar way, if we restrict "\Delta z" to take only imaginary values (in this case then "\Delta z=i\Delta y"), then we conclude that there exists a partial derivative "\partial f/\partial y" at the point "z":

"f_y(z):=\frac{\partial f}{\partial y}(z):=\lim\limits_{\Delta y\to 0}\frac{f(z+i\Delta y)-f(z)}{\Delta y}=i\lim\limits_{\Delta y\to 0}\frac{f(z+i\Delta y)-f(z)}{i\Delta y}=if'(z)"

Let "u(z)" and "v(z)" are (respectively) the real and the imaginary part of the function "f(z)". Then

"f(z)=u(z)+iv(z)",

"f_x(z)=u_x(z)+iv_x(z)" and

"u_x(z)^2+v_x(z)^2=|f_x(z)|^2=|f'(z)|^2"

Similarly,

"f_y(z)=u_y(z)+iv_y(z)" and

"u_y(z)^2+v_y(z)^2=|f_y(z)|^2=|if'(z)|^2=|f'(z)|^2"

Therefore, "|f'(z)|^2=u_x(z)^2+v_x(z)^2=u_y(z)^2+v_y(z)^2", as required.